OpenStudy (anonymous):

Will medal! Please help!! How many solutions does this system of equations have? y = –3x + 7 y = –3x – 6

OpenStudy (campbell_st):

well do you notice anything about the slopes of the lines..?

OpenStudy (anonymous):

How do I know the slopes of the lines? @campbell_st

OpenStudy (campbell_st):

ok... the basic slope intercept form of a line is \[y = mx + b\] m is the slope of the line and b is where it cuts the y -axis so any comment about the slope

OpenStudy (anonymous):

Nope because I'm completely lost. I truthfully suck at math so... @campbell_st

OpenStudy (campbell_st):

ok... here is an easy solution graph them using https://www.desmos.com/calculator on the left side just type the equation, press enter then type the 2nd equation. see what happens

OpenStudy (anonymous):

Do you have answer choices to this question too?

OpenStudy (anonymous):

Yes I do @Najia2000 & I typed it in campbell_st

OpenStudy (campbell_st):

can we ignore the answers for a while and concentrate on improving understanding

OpenStudy (anonymous):

Sure, of course

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@Najia2000 1 infinitely many 0 2

OpenStudy (campbell_st):

so use the link to graph the lines...

OpenStudy (anonymous):

Okay but why do I need to

OpenStudy (anonymous):

Is there only one solution?

OpenStudy (campbell_st):

it will given you a graph of what the lines look like...

OpenStudy (campbell_st):

why guess an answer..

OpenStudy (anonymous):

I'm not guessing, I'm just wondering. I literally have no idea how this works, you're not really explaining it.

OpenStudy (anonymous):

Actually, it's no solution isn't it? Because the lines parallel?

OpenStudy (campbell_st):

ok... if you attempted to graph them you would find they look like this

OpenStudy (campbell_st):

now I asked you about the slopes...at the beginning the general slope intercept form is y = mx + b your lines y = -3x + 7 y = -3x - 6 so any thoughts on what the slope of each line is..?

OpenStudy (anonymous):

I did attempt to graph it, and that is what I got. Please explain further.

OpenStudy (campbell_st):

can you make any general comment about the lines you graphed..?

OpenStudy (anonymous):

They don't touch, that's all I know.

OpenStudy (campbell_st):

your teacher would have talked about parallel lines in class... as the have the same slope.... -3 the answer choices are asking how many times the lines intersect, or cross... this is called the solution to the system of equations... so any thoughts on how many times they intersect..?

OpenStudy (anonymous):

Literally none. So I was right then?

OpenStudy (campbell_st):

that's correct... so if you can't identify the slopes... then use something to graph them... the image helps your understanding